OpenStudy (anonymous):

5. Solve the inequality. Show your work. 6y – 8 ≤ 10

OpenStudy (michele_laino):

please add 8 to both sides of your inequality, first

OpenStudy (anonymous):

\[6y-8\le10\rightarrow 6y \le8+10 \rightarrow6y \le 18 \]

OpenStudy (anonymous):

\[y \le \frac{ 18 }{ 6 } \rightarrow y \le 3\]

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

thank you